OpenStudy (anonymous):

Prove: (can only solve from left side to prove right side) (tanx+tany)/(cotx+coty)=tanx*tany

OpenStudy (loser66):

denominator : cot x + cot y = \(\dfrac{1}{tanx}+\dfrac{1}{tany}\) = \(\dfrac{tany +tanx}{tanxtany}\) now, put them under the numerator